# Latent Adversarial Training Improves Robustness to Persistent Harmful Behaviors in LLMs

### v2

@@ Abstract
Large language models (LLMs) can often be made to behave in undesirable ways that they are explicitly fine-tuned not to. For example, the LLM red-teaming literature has produced a wide variety of ‘jailbreaking’ techniques to elicit harmful text from models that were fine-tuned to be harmless. Recent work on red-teaming, model editing, and interpretability suggests that this challenge stems from how (adversarial) fine-tuning largely serves to suppress rather than remove undesirable capabilities from LLMs. Prior work has introduced latent adversarial training (LAT) as a way to improve robustness to broad classes of failures. These prior works have considered untargeted latent space attacks where the adversary perturbs latent activations to maximize loss on examples of desirable behavior. Untargeted LAT can provide a generic type of robustness but does not leverage information about specific failure modes. Here, we experiment with targeted LAT where the adversary seeks to minimize loss on a specific competing task. We find that it can augment a wide variety of state-of-the-art methods. First, we use targeted LAT to improve robustness to jailbreaks, outperforming a strong R2D2 baseline with orders of magnitude less compute. Second, we use it to more effectively remove backdoors with no knowledge of the trigger. Finally, we use it to more effectively unlearn knowledge for specific undesirable tasks in a way that is also more robust to re-learning. Overall, our results suggest that targeted LAT can be an effective tool for defending against harmful behaviors from LLMs.111Code is available at github.com/aengusl/latent-adversarial-training. Models are available at huggingface.co/LLM-LAT. Chat with our jailbreaking robust model at abhayesian.com/lat-chat..
@@ 1 Introduction
Despite efforts from developers to remove harmful capabilities from large language models (LLMs), they can persistently exhibit undesirable behaviors. For example, a string of recent red-teaming papers [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have demonstrated diverse techniques that can be used to elicit instructions for building bombs from state-of-the-art LLMs. Developers have made progress on these problems using improved data (e.g., [19]) and adversarial training (e.g., [20, 21]). However, some harmful capabilities resist removal via fine-tuning and have been a persistent challenge toward building more trustworthy models [22, 23].
Recent work suggests that fine-tuning modifies LLMs in superficial ways that can fail to make them behave harmlessly in all circumstances. Research on interpretability [24, 25, 26, 27, 28, 29], representation engineering [30, 31, 32], continual learning [33, 34, 35, 36, 37, 38, 39, 40], and fine-tuning [25, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50] has suggested that fine-tuning struggles to make fundamental changes to an LLM’s inner knowledge and capabilities. For example, Jain et al. [25] likened fine-tuning in LLMs to merely modifying a “wrapper” around a stable, general-purpose set of latent capabilities.
In this paper, we use latent adversarial training (LAT) [51, 52] to make LLMs more robust to exhibiting persistent unwanted behaviors. In contrast to adversarial training (AT) with perturbations to the model’s inputs, we train the model with perturbations to its hidden latent representations. Because models represent features at a higher level of abstraction in the latent space [53], we hypothesize that LAT can better facilitate the removal of neural circuitry responsible for unwanted behaviors. Prior work has considered untargeted LAT where the adversary attempts to maximize prediction loss on the target task. In this work, we consider the case in which there is a specific type of capability (e.g., a backdoor) that we want to remove. Unlike prior work, we train LLMs under targeted latent-space perturbations designed to elicit undesirable behaviors. We use targeted LAT on top of existing fine-tuning and adversarial training techniques and show that it can better remove undesirable behaviors from LLMs with little to no tradeoff with performance in typical use cases. We make two contributions:
@@
We propose targeted latent adversarial training (LAT) as a way to more thoroughly remove undesirable behaviors from LLMs.
@@
We show that targeted LAT can combine with and improve over a wide range of state-of-the-art techniques.
In Section 4.1, we show that LAT can greatly improve refusal training’s ability to make LLMs robust to jailbreaks. We find that LAT outperforms R2D2 [21] with orders of magnitude less compute.
In Section 4.2, we use LAT to greatly improve DPO’s [54] ability to remove LLM backdoors when the trigger is unknown and the response is only vaguely specified. Our results suggest that LAT is a solution to the ‘Sleeper Agent’ problem posed in Hubinger et al. [55].
In Section 4.3, we use LAT to improve on the abilities of WHP [56], gradient ascent [57], and RMU [58] to unlearn unwanted knowledge. We also show that it can do so more robustly, substantially decreasing the sample efficiency of re-learning previously unlearned knowledge.
@@ 2 Related Work
Latent-space attacks and LAT have been previously studied in vision models [51, 59, 60, 61, 62] and language models [31, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74]. Our work is closely related to Casper et al. [52], who used LAT to defend against backdoors and unforeseen classes of adversarial attacks. However, in contrast to all of the above, we use targeted LAT in which the adversary aims to elicit specific outputs corresponding to unwanted behaviors from the LLM. This makes our work similar to concurrent work by Xhonneux et al. [75] who perform targeted adversarial training, but in the model’s embedding space instead of its latents and Zeng et al. [76] who perform targeted LAT, but only for the task of backdoor removal. Meanwhile, several works have shown that the high-level behaviors of LLMs can be altered using perturbations to their internal activations [2, 77, 78, 79, 80, 81, 82, 83, 84, 85], but, to the best of our knowledge, these perturbations have not been trained against to improve robustness. Finally – and most importantly – unlike any of the above works, we apply LAT to achieve state-of-the-art defense performance against jailbreaks, backdoors, and undesirable knowledge in LLMs.
@@
Multiple techniques have been used to make LLMs behave more robustly including data preprocessing [86, 87, 88], scaling [89, 90],222Although increasing scale can also exacerbate some vulnerabilities [91, 92, 93, 94]). and adversarial training (AT) [20, 95, 96, 97, 98]. However, state-of-the-art LLMs persistently display vulnerabilities to novel attacks [11, 99, 100]. Meanwhile, Hubinger et al. [55], Jain et al. [86], Pawelczyk et al. [101], and Casper et al. [52] have each shown cases in which AT can fail to fix specific problems with LLMs that occur off the attack distribution used for training. In this paper, we demonstrate that robustness to unseen jailbreak and backdoor attacks can be improved using LAT.
@@
Large language models are vulnerable to threats from backdoors (also known as trojans). Typically, these threats arise from a malicious actor poisoning pretraining or fine-tuning data to make the model exhibit harmful behaviors upon encountering some arbitrary trigger [102, 103, 104, 105, 106, 107]. One motivation for studying LLM backdoors is the practical threat they pose [106]. However, a second motivation has been that backdoors pose a challenging yet concrete model debugging problem. Addressing backdoors is difficult because, without knowledge of the trigger (which can be an arbitrary input feature), it is difficult to train the model in a way that removes the backdoor. Hubinger et al. [55] found that adversarial training could even strengthen a “sleeper agent” backdoor, which they designed to make the model behave harmfully upon seeing prompts indicating that the year was 2024.
@@
Machine unlearning has historically been motivated by erasing the influence of data on a trained model to reduce risks to privacy or violations of fair use. However, in LLMs, unlearning is increasingly motivated by removing harmful capabilities of models [108, 58]. Prior works have introduced a number of LLM unlearning techniques [28, 56, 57, 58, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125]. Here, we show that LAT can improve over unlearning techniques including state-of-the-art RMU [58].
@@ 3 Methods
We can view an LLM with parameters , as a composition of two functions, , where is a feature extractor which maps text to latent activations and maps those latent activations to output a probability distribution for sampling: i.e., . We define an adversarial attack as a function with parameters which modifies the LLM’s inputs or latent activations.
During standard AT, the model is trained to be robust to attacks in the input space via some training loss function, . The training objective is thus . In contrast, during latent adversarial training (LAT), the model is instead trained to be robust to attacks to the latent activations:
During untargeted LAT (e.g., [52]), the attacker seeks to steer the model away from the desired behavior on a training example . The attacker’s objective is thus . However, during targeted LAT, the attacker seeks to steer the model toward some undesirable target behavior :
@@
Performing basic targeted LAT requires a dataset of desirable behaviors and a dataset of undesirable behaviors . For us, in most cases, this takes the form of prompts and paired harmless and harmful completions . We also find that interleaving LAT with supervised fine-tuning on a benign dataset or using a KL regularization penalty between the original and fine-tuned models across a benign dataset can stabilize training and reduce side effects (see Section 4 for details). We refer to this benign dataset as . Here, as done in Casper et al. [52], we attack the residual stream of transformer LLMs with -norm-bounded perturbations, calculated using projected gradient descent (PGD) [126]. Because the model and attacker are optimized using different completions to prompts, we only perturb the token positions in the residual stream corresponding to the prompt (and not the completions) – see Figure 1. In practice, we found that perturbing the residual stream at multiple layers rather than a single layer, each with its own constraint typically yielded better results. After experimenting with different choices of layers, we converged toward the heuristic of perturbing four layers, evenly spaced throughout the network. In all experiments, we performed hyperparameter sweeps to select a perturbation bound.
@@ 4 Experiments
| Jailbreak Robustness (Section 4.1) | Refusal Training (RT) |
| Embedding-Space Adversarial Training [75] | |
| Backdoor Removal (Section 4.2) | Direct Preference Optimization (DPO) [54] |
| Unlearning (Section 4.3) | Who’s Harry Potter (WHP) [56] |
| Gradient Ascent (GA) [57] | |
| Representation Misdirection for Unlearning (RMU) [58] |
@@
Here, we experiment with targeted LAT for improving robustness to jailbreaks, unlearning undesirable knowledge, and removing backdoors. Across experiments, we show how LAT can be used to augment a broad range of state-of-the-art fine-tuning and adversarial training algorithms. Table 1 summarizes the methods we augment with targeted LAT.333All experiments presented below were run on a single A100 or H100 GPU except for ones involving R2D2 [58] in Section 4.1 which were run on eight. All model training runs lasted less than 12 hours of wall-clock time.
@@ 4.1 Improving Robustness to Jailbreaks
Here, we demonstrate that targeted LAT can be helpful for making models more resistant to exhibiting unwanted behaviors via jailbreaking attacks with minimal side effects.
@@
We create a dataset of triples containing: prompts, harmful completions, and harmless completions using a method based on Self-Instruct [127]. We first generate a set of harmful user requests by few-shot prompting Mistral-7B [128] with harmful requests seeded by AdvBench [129]. We then filter for prompts of an intermediate length and subsample for diversity by clustering BERT embeddings [130] and sampling one prompt from each cluster. To generate harmful responses to the harmful user requests, we sampled from Zephyr-7B-Beta which was fine-tuned from Mistral-7B [128] by Tunstall et al. [131] to respond helpfully to user requests. We similarly generate refusals (harmless responses) using Llama2-7B-chat [96] instruction-prompted to refuse harmful requests.
@@
We fine-tune Llama2-7B-chat [96] using refusal training (RT). We implement refusal training based on Mazeika et al. [21] using both a ‘toward’ and ‘away’ loss term calculated with respect to harmless/harmful example pairs. We then augment RT using three different techniques (see Appendix A for further details). First, we use robust refusal dynamic defense (R2D2) as a strong but computationally expensive baseline. R2D2 is an adversarial training technique based on synthesizing adversarial suffixes using greedy coordinate gradient (GCG) attacks [129].444We also experimented with R2D2-LAT but found it to result in unstable training. We leave further experimentation with R2D2-LAT to future work. Second, we augment RT using embedding-space adversarial training (RT-EAT) [75]. We refer to this as RT-EAT. Finally, we augment RT-EAT using LAT (RT-EAT-LAT). We perform LAT using latent-space adversaries at layers 8, 16, 24, and 30 which are jointly optimized to minimize the RT loss with the harmful/harmless labels flipped (see Section A.1). Additionally, we also experiment with Llama3-8B [132]. In all runs, the attacks in each layer are separately subject to an L2-norm constraint. In all experiments, we use the UltraChat dataset [133] as a benign fine-tuning dataset to preserve the model’s performance. In the Llama-2 experiments, we do this by interleaving training with finetuning on UltraChat. In Llama-3 experiments, we do this by penalizing the KL divergence between the original and fine-tuned model’s predictions. Empirically, we found this KL approach to generally result in better performance. Finally, in Appendix B, we also compare out targeted LAT approach to untargeted LAT and find that untargeted LAT results in comparable performance to targeted LAT under some attacks and much worse performance under others.
@@
To evaluate the models’ performance in non-adversarial settings, we use the Massive Multitask Language Understanding (MMLU) benchmark, [134], the MT-Bench benchmark (using a single-turn version) [135], and the models’ rate of compliance with benign requests. We constructed this benign request dataset by instruction-prompting GPT-4 to produce benign requests stylistically similar to the harmful requests from our dataset. Similar to Liu et al. [7], we count refusals based on string-matching refusal phrases (this was only done to calculate the “Compliance” column of Table 2). Next, to measure robustness, we use six attacks: direct requests with no adversarial optimization, prefilling attacks [136], PAIR [9], AutoPrompt (AP) attacks [137], greedy coordinate gradient (GCG) attacks [129], and many-shot jailbreaking attacks [18] combined with GCG. We evaluate the success of attacks using the StrongReject autograder [138], a GPT-4o based autograder designed to classify successful jailbreak attempts.555The StrongReject autograder [138] was designed to minimize the rate at which unsuccessful attacks were mistakenly classified as successful. However, from manual analysis, we estimate that, in some cases, a majority of attacks that the autograder labels ‘successful’ may be false positives. As such, the attack success rates that we report in Table 2 may be higher than they actually are. In Appendix C, we also report results for the HarmBench autograder [21] and find similar overall trends. Finally, we estimate compute as in Xhonneux et al. [75] by calculating the total number of forward and backward passes used during training, ignoring batch and device parallelism.
Targeted LAT improves robustness to jailbreaks with minimal side effects.
Table 2 presents results. Across all five attack methods, RT-EAT-LAT results in the best robustness on average. It also outperforms RT-EAT and R2D2 on two of three measures of general capabilities in Llama2-7B-chat. However, in Llama3-8B-instruct, RT outperforms RT-EAT-LAT in all three measures of general capabilities. Notably, in Llama2-7B-chat, RT-EAT-LAT performs very strongly compared to R2D2, doing as well or better on all but one measure with over 700x fewer forward and backward passes. Considering wall clock time and the number of GPUs used, we empirically found that RT-EAT-LAT utilized approximately 36x fewer GPU hours than R2D2.

### v3

@@ Abstract
Large language models (LLMs) can often be made to behave in undesirable ways that they are explicitly fine-tuned not to. For example, the LLM red-teaming literature has produced a wide variety of ‘jailbreaking’ techniques to elicit harmful text from models that were fine-tuned to be harmless. Recent work on red-teaming, model editing, and interpretability suggests that this challenge stems from how (adversarial) fine-tuning largely serves to suppress rather than remove undesirable capabilities from LLMs. Prior work has introduced latent adversarial training (LAT) as a way to improve robustness to broad classes of failures. These prior works have considered untargeted latent space attacks where the adversary perturbs latent activations to maximize loss on examples of desirable behavior. Untargeted LAT can provide a generic type of robustness but does not leverage information about specific failure modes. Here, we experiment with targeted LAT where the adversary seeks to minimize loss on a specific competing task. We find that it can augment a wide variety of state-of-the-art methods. First, we use targeted LAT to improve robustness to jailbreaks, outperforming a strong R2D2 baseline with orders of magnitude less compute. Second, we use it to more effectively remove backdoors with no knowledge of the trigger. Finally, we use it to more effectively unlearn knowledge for specific undesirable tasks in a way that is also more robust to re-learning. Overall, our results suggest that targeted LAT can be an effective tool for defending against harmful behaviors from LLMs. 111Code is available at github.com/aengusl/latent-adversarial-training. Models are available at huggingface.co/LLM-LAT. Chat with our jailbreaking robust model at abhayesian.com/lat-chat.
@@ 1 Introduction
Despite efforts from developers to remove harmful capabilities from large language models (LLMs), they can persistently exhibit undesirable behaviors. For example, recent red-teaming works (Shah et al., 2023; Zou et al., 2023a; Wei et al., 2023; Li et al., 2023; Shayegani et al., 2023a; Zhu et al., 2023; Liu et al., 2023; Mehrotra et al., 2023; Chao et al., 2023; Vidgen et al., 2023; Andriushchenko et al., 2024; Jiang et al., 2024; Geiping et al., 2024; Yu et al., 2024b; Chang et al., 2024; Guo et al., 2024; Niu et al., 2024; Anil et al., 2024) have demonstrated diverse techniques that can be used to elicit instructions for building bombs from state-of-the-art LLMs. Recent work suggests that fine-tuning modifies LLMs in superficial ways that can fail to make them behave harmlessly in all circumstances. Research on interpretability (Juneja et al., 2022; Jain et al., 2023b; Lubana et al., 2023; Prakash et al., 2024; Patil et al., 2023; Lee et al., 2024), representation engineering (Wei et al., 2024; Schwinn et al., 2024; Li et al., 2024b), continual learning (Ramasesh et al., 2021; Cossu et al., 2022; Li et al., 2022; Scialom et al., 2022; Luo et al., 2023; Kotha et al., 2023; Shi et al., 2023; Schwarzschild et al., 2024), and fine-tuning (Jain et al., 2023b; Yang et al., 2023; Qi et al., 2023; Bhardwaj & Poria, 2023; Lermen et al., 2023; Zhan et al., 2023; Ji et al., 2024; Qi et al., 2024; Hu et al., 2024; Halawi et al., ; Greenblatt et al., 2024; Deeb & Roger, 2024) has suggested that fine-tuning struggles to make fundamental changes to an LLM’s inner knowledge and capabilities.
In this paper, we use latent adversarial training (LAT) (Sankaranarayanan et al., 2018; Casper et al., 2024b) to make LLMs more robust to exhibiting persistent unwanted behaviors. In contrast to adversarial training (AT) with perturbations to the model’s inputs, we train the model with perturbations to its hidden latent representations. Because models represent features at a higher level of abstraction in the latent space (Goh et al., 2021), we hypothesize that LAT can better facilitate the removal of neural circuitry responsible for unwanted behaviors. Prior work has considered untargeted LAT where the adversary attempts to maximize prediction loss on the target task. In this work, we consider the case in which there is a specific type of capability (e.g., a backdoor) that we want to remove. Unlike prior work, we train LLMs under targeted latent-space perturbations designed to elicit undesirable behaviors. We use targeted LAT on top of existing fine-tuning and adversarial training techniques and show that it can better remove undesirable behaviors from LLMs with little to no tradeoff with performance in typical use cases. We make two contributions:
@@
We propose targeted latent adversarial training (LAT) as a way to more thoroughly remove persistent undesirable behaviors from LLMs.
@@
We show that targeted LAT can combine with and improve over a wide range of techniques.
In Section˜4.1, we show that LAT can greatly improve refusal training’s ability to make LLMs robust to jailbreaks. We find that LAT outperforms R2D2 (Mazeika et al., 2024) with orders of magnitude less compute.
In Section˜4.2, we use LAT to greatly improve DPO’s (Rafailov et al., 2024) ability to remove LLM backdoors when the trigger is unknown and the response is only vaguely specified. Our results suggest that LAT is a solution to the ‘Sleeper Agent’ problem posed in Hubinger et al. (2024).
In Section˜4.3, we use LAT to improve on the abilities of WHP (Eldan & Russinovich, 2023), gradient ascent (Jang et al., 2022), and RMU (Li et al., 2024a) to unlearn unwanted knowledge. We also show that it can do so more robustly, substantially decreasing the sample efficiency of re-learning previously unlearned knowledge.
@@ 2 Related Work
Latent-space adversarial modifications and LAT have been previously studied in vision models (Sankaranarayanan et al., 2018; Singh et al., 2019; Park & Lee, 2021; Qian et al., 2021; Zhang et al., 2023b) and language models (Schwinn et al., 2024; Jiang et al., 2019; Zhu et al., 2019; Liu et al., 2020; He et al., 2020; Kuang & Bharti, ; Li & Qiu, 2021; Sae-Lim & Phoomvuthisarn, 2022; Pan et al., 2022; Schwinn et al., 2023; Geisler et al., 2024; Fort, 2023; Kitada & Iyatomi, 2023). Our work is closely related to Casper et al. (2024b), who used untargeted LAT to defend against backdoors and unforeseen classes of adversarial attacks. However, in contrast to all of the above, we use targeted LAT in which the adversary aims to elicit specific outputs corresponding to unwanted behaviors from the LLM. See also concurrent work by Xhonneux et al. (2024) who perform AT on the model’s text embeddings, Zeng et al. (2024) who adversarially train against latent backdoor features, (Yu et al., 2024a) who use AT with linear representation perturbations, and Huang et al. (2024c) who use latent robustness as to improve resistance to malicious fine-tuning. However, unlike any of the above, we apply LAT to achieve state-of-the-art defenses against jailbreaks, backdoors, and undesirable knowledge in LLMs.
@@
Multiple techniques have been used to make LLMs behave more robustly including adversarial training (AT) (Ziegler et al., 2022; Ganguli et al., 2022; Touvron et al., 2023; Achiam et al., 2023; Team et al., 2023). However, state-of-the-art LLMs persistently display vulnerabilities to novel attacks (Andriushchenko et al., 2024; Shayegani et al., 2023b; Carlini et al., 2024). Meanwhile, Hubinger et al. (2024), Jain et al. (2023a), Pawelczyk et al. (2024), and Casper et al. (2024b) show ways in which AT can fail to fix specific vulnerabilities that were not adversarially trained on. Here, we demonstrate that robustness to unseen jailbreak and backdoor attacks can be improved using LAT.
@@
Large language models are vulnerable to threats from backdoors (also known as trojans). Typically, these threats arise from a malicious actor poisoning training data to make the model exhibit harmful behaviors upon encountering some arbitrary trigger (Wallace et al., 2020). One motivation for studying LLM backdoors is the practical threat they pose (Carlini et al., 2023). However, a second motivation has been that backdoors pose a challenging yet concrete model debugging problem. Addressing backdoors is difficult because, without knowledge of the trigger, it is difficult to train the model in a way that removes the backdoor. Hubinger et al. (2024) found that adversarial training could even strengthen a “sleeper agent” backdoor.
@@
In LLMs, machine unlearning is increasingly motivated by removing harmful capabilities of models (Liu et al., 2024a; Li et al., 2024a). Prior works have introduced a number of LLM unlearning techniques (Eldan & Russinovich, 2023; Li et al., 2024a; Lu et al., 2022; Yao et al., 2023; Chen & Yang, 2023; Ishibashi & Shimodaira, 2023; Yu et al., 2023; Wang et al., 2023; Wu et al., 2023; Zhang et al., 2023a; Yuan et al., 2023; Maini et al., 2024; Lu et al., 2024; Goel et al., 2022; Lo et al., 2024; Huang et al., 2024a; Liu et al., 2024b), but existing methods suffer from adversarial vulnerabilities (Lynch et al., 2024; Łucki et al., 2024). Here, we show that LAT can improve over unlearning techniques including state-of-the-art RMU (Li et al., 2024a).
@@ 3 Methods
We can view an LLM with parameters , as a composition of two functions, , where is a feature extractor which maps text to latent activations and maps those latent activations to output a probability distribution for sampling: i.e., . We define an adversarial attack as a function with parameters which modifies the LLM’s inputs or latent activations. During standard AT, the model is trained to be robust to attacks in the input space via some training loss function, . The training objective is thus . In contrast, during latent adversarial training (LAT), the model is instead trained to be robust to attacks to the latent activations:
During untargeted LAT (e.g., Casper et al. (2024b)), the attacker seeks to steer the model away from the desired behavior on a training example . The attacker’s objective is thus . However, during targeted LAT, the attacker seeks to steer the model toward some undesirable target behavior :
@@
Performing basic targeted LAT requires a dataset of desirable behaviors and a dataset of undesirable behaviors . For us, in most cases, this takes the form of prompts and paired harmless and harmful completions . We also find that interleaving LAT with supervised fine-tuning on a benign dataset or using a KL regularization penalty between the original and fine-tuned models across a benign dataset can stabilize training and reduce side effects (see Section˜4 for details). We refer to this benign dataset as . We attack the residual stream of transformer LLMs with -norm-bounded perturbations, calculated using projected gradient descent (PGD) (Madry et al., 2017). Because the model and attacker are optimized using different completions to prompts, we only perturb the positions in the residual stream corresponding to the prompt – see Figure˜1. We found that perturbing the residual stream at multiple layers rather than a single layer, each with its own constraint typically yielded better results. After experimenting with different choices of layers (1, 2, 3, 4, 10, 16, 22, and 28), we found that the simple heuristic of selecting four evenly spaced layers worked well across models and experiments. We empirically selected the perturbation bound through a grid search over , choosing the value that provided maximal robustness against jailbreak attacks on Llama-2. Notably, these hyperparameters demonstrated good generalization, maintaining their effectiveness when held fixed across our subsequent experiments on different models and tasks.
@@ 4 Experiments
| Jailbreak Robustness (Section˜4.1) | Refusal Training (RT) |
| Embedding-Space Adversarial Training (Xhonneux et al., 2024) | |
| Backdoor Removal (Section˜4.2) | Direct Preference Optimization (DPO) (Rafailov et al., 2024) |
| Unlearning (Section˜4.3) | Who’s Harry Potter (WHP) (Eldan & Russinovich, 2023) |
| Gradient Ascent (GA) (Jang et al., 2022) | |
| Representation Misdirection for Unlearning (RMU) (Li et al., 2024a) |
@@
Here, we experiment with targeted LAT for improving robustness to jailbreaks, unlearning undesirable knowledge, and removing backdoors. Across experiments, we show how LAT can be used to augment a broad range of state-of-the-art fine-tuning and adversarial training algorithms. Table˜1 summarizes the methods we augment with targeted LAT.222All experiments were run on a single A100 or H100 GPU except for ones involving R2D2 (Li et al., 2024a) in Section 4.1 which were run on eight. All training runs lasted less than 12 hours of wall-clock time.
@@ 4.1 Improving Robustness to Jailbreaks
We create a dataset of triples containing: prompts, harmful completions, and harmless completions using a method based on Self-Instruct (Wang et al., 2022). We first generate a set of harmful user requests by few-shot prompting Mistral-7B (Jiang et al., 2023) with harmful requests seeded by AdvBench (Zou et al., 2023b). We then filter for prompts of an intermediate length and subsample for diversity by clustering BERT embeddings (Devlin et al., 2018) and sampling one prompt from each cluster. To generate harmful responses to the harmful user requests, we sampled from Zephyr-7B-Beta which was fine-tuned from Mistral-7B (Jiang et al., 2023) by Tunstall et al. (2023) to respond helpfully to user requests. We similarly generate refusals (harmless responses) using Llama2-7B-chat (Touvron et al., 2023) instruction-prompted to refuse harmful requests.
@@
Here, we fine-tune models using refusal training (RT). We implement refusal training based on Mazeika et al. (2024) using both a ‘toward’ and ‘away’ loss term calculated with respect to harmless/harmful example pairs. We then augment RT using three different techniques (see Appendix B for further details). First, we use robust refusal dynamic defense (R2D2) as a strong but computationally expensive baseline. Second, we augment RT using embedding-space (i.e. latent layer zero) adversarial training (RT-EAT) (Xhonneux et al., 2024). We refer to this as RT-EAT. Finally, we augment RT-EAT using LAT (RT-EAT-LAT). We perform LAT using latent-space adversaries at layers 8, 16, 24, and 30 which are jointly optimized to minimize the RT loss with the harmful/harmless labels flipped (see Section˜B.1). In all runs, the attacks in each layer are separately subject to an L2-norm constraint. In all experiments, we use the UltraChat dataset (Ding et al., 2023) as a benign fine-tuning dataset to preserve the model’s performance. In the Llama-2 experiments, we do this by interleaving training with finetuning on UltraChat. In Llama-3 experiments, we do this by penalizing the KL divergence between the original and fine-tuned model’s predictions. Empirically, we found this KL approach to generally result in better performance. Finally, in Appendix˜C, we also compare oue targeted LAT approach to untargeted LAT and find that untargeted LAT results in comparable performance to targeted LAT under some attacks and much worse performance under others.
@@
To evaluate the models’ performance in non-adversarial settings, we use the Massive Multitask Language Understanding (MMLU) benchmark, (Hendrycks et al., 2020), the MT-Bench benchmark (using a single-turn version) (Zheng et al., 2024), and the models’ rate of compliance with benign requests. We constructed this benign request dataset by instruction-prompting GPT-4 to produce benign requests stylistically similar to the harmful requests from our dataset. Similar to Liu et al. (2023), we count refusals based on string-matching refusal phrases (this was only done to calculate the “Compliance” column of Table˜2). Next, to measure robustness, we use six attacks: direct requests with no adversarial optimization, prefilling attacks (Haizelabs, ), PAIR (Chao et al., 2023), AutoPrompt (AP) attacks (Shin et al., 2020), greedy coordinate gradient (GCG) attacks (Zou et al., 2023b), and many-shot jailbreaking attacks (Anil et al., 2024) combined with GCG. We evaluate the success of attacks using the StrongReject autograder (Souly et al., 2024), a GPT-4o based autograder designed to classify successful jailbreak attempts.333The StrongReject autograder (Souly et al., 2024) was designed to minimize the rate at which unsuccessful attacks were mistakenly classified as successful. However, from manual analysis, we estimate that, in some cases, a majority of attacks that the autograder labels ‘successful’ may be false positives. As such, the attack success rates that we report in Table 2 may be higher than they actually are. In Appendix˜D, we also report results for the HarmBench autograder (Mazeika et al., 2024) and find similar overall trends. Finally, we estimate compute as in Xhonneux et al. (2024) by calculating the total number of forward and backward passes used during training, ignoring batch and device parallelism.
LAT improves robustness to jailbreaks with minimal side effects.
Table˜2 presents results. Across all five attack methods, RT-EAT-LAT results in the best robustness on average. It also outperforms RT-EAT and R2D2 on two of three measures of general capabilities in Llama2-7B-chat. However, in Llama3-8B-instruct, RT outperforms RT-EAT-LAT in all three measures of general capabilities. Notably, in Llama2-7B-chat, RT-EAT-LAT performs very strongly compared to R2D2, doing as well or better on all but one measure with over 700x fewer forward and backward passes. Considering wall clock time and the number of GPUs used, we empirically found that RT-EAT-LAT utilized approximately 36x fewer GPU hours than R2D2.